When heated, KClO3 decomposes into KCl and O2.

2 KClO3 ⟶ 2 KCl + 3 O2

If this reaction produced 53.7 g KCl, how many grams of O2 were produced?
1
Expert's answer
2019-10-08T07:40:22-0400

The molar mass of potassium chloride is 39+35.5 = 74.5. Thus formed 53.7/74.5 = 0.72 mol oxygen. Based on the oxygen reaction equation formed (0.72*2)/3 = 0.48 mol or 0.48*22.4 = 10.75 liters